how do you write y=5-3x in standard forum
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)


Hi
standard form of a line is in the form Ax + By = C where A >0
y = 5 - 3x
3x + y = 5 In standard form

Answer by stanbon(75887)

write y=5-3x in standard forum
----
Standard form is ax+by=c
Your Problem:
3x+y = 5
--------------
Cheers,
Stan H.